OpenStudy (anonymous):

Determine the solution set of (x + 2)^2 = 36. {4, -8}?

OpenStudy (kainui):

So when you take the square root of both sides you can either use positive or negative 6 to get 36, right? So this is where you have two separate answers come from!

OpenStudy (kainui):

Just write them out as two separate equations, x+2=6 and x+2=-6 and keep solving, you see?

OpenStudy (anonymous):

{6, -8}?

OpenStudy (kainui):

No, it's {4,-8} because you are still subtracting 2 from each side to solve for x. So 6-2=4 and -6-2=8. Only the 6 is changing its sign.

OpenStudy (anonymous):

oh, okay thank you!